OpenStudy (anonymous):

Which is a solution to the system of equations below? 7x = 2y + 3 y = 3x + 1

OpenStudy (anonymous):

7x-2y=3 -3x+y=1 7x-2y=3 (2)(-3x+y=1) 7x-2y=3 -6x+2y=2 x=5 y=16

OpenStudy (salina):

7x = 2(3x + 1) + 3 7x = 6x+2 +3 7x-6x=5 x=5 y = 3(5) + 1 y= 16

OpenStudy (anonymous):

thanks!